# A non-abelian version of Deligne’s Fixed Part Theorem

### v3

@@ 1. Introduction
If we furthermore assume $f$ to be proper then these conditions are also equivalent to:
@@
    the filtration ${\mathsf{F}}_{s}$ extends to a (relative) Griffiths transversal filtration by subbundles $\hat{\mathsf{F}}$ of the formal isomonodromic extension (^,^:^$(\hat{\mathcal{E}},\hat{\nabla}\colon\hat{\mathcal{E}}\to\hat{\Omega}^{1}_{\hat{X}/\hat{S}}(\hat{\mathcal{E}}))$.
Here “essentially extends” means that it extends after base change with respect to a finite, étale, surjective covering of $S$, see Subsection 3.1 for a precise formulation. In 5) $\hat{S}$ is the formal scheme of $S$ along $s$ and $\hat{X}$ the formal scheme of $X$ along $X_{s}$. The formal isomonodromic deformation $(\hat{\mathcal{E}},\hat{\nabla})$ is the formal completion of the flat bundle on $X_{\Delta}$ associated to the local system on $X_{\Delta}$ which canonically extends $\mathbb{L}_{s}$.
@@
Acknowledgment. It is a pleasure to thank Alexander Petrov for numerous enlightening discussions and for sharing his insights into isomonodromic deformations of Hodge structures. We thank Claude Sabbah for kindly helping us with references on variations of Hodge structure. We thank Daniel Litt for various helpful discussions on local systems in particular on the condition 5) in Theorem 1.1. We are grateful to Takuro Mochizuki for explaining to us what is known and what can be expected about families of pluri-harmonic metrics. We thank Xiwen Zhu for an enlightening discussion on the analogy between Theorem 1.1 and his own theorem on the preservation of the de Rham property of a $p$-adic local system, see [LZ17].
@@ 2.2. Geometric examples
Let $X$ and $S$ be connected complex manifolds and let $f\colon X\to S$ be a good holomorphic map as defined in Section 1. Fix a point $x\in X$ and set $s=f(x)$. By Thom’s first isotopy lemma [GM88, Thm. I.1.5] or Ehresmann’s theorem in the proper case, $f$ is a topological fibration. We obtain an exact sequence of homotopy groups
@@ 3.1. Integral Fixed Part Theorem
As $f$ is a topological fibration, $\pi_{1}(S,s)$ acts on the cohomology $H^{i}(X_{s},{\mathbb{Q}})$. We denote by $\widehat{S}$ the formal completion of $S$ along $s$ and by $\widehat{X}$ the formal completion of $X$ along $X_{s}$.
@@ 5.1. Polarized variations of ${\mathbb{C}}$-Hodge structure
Let us recall for later reference one of the main results about degeneration of pure ${\mathbb{C}}$-Hodge structure. Let $\mathbb{L}$ be a unipotent complex local system on $\Delta^{\times}=\Delta\setminus\{0\}$ with $\Delta=\{z\in{\mathbb{C}}\,|\,|z|<1\}$. Let $Q\colon\mathbb{L}\times\overline{\mathbb{L}}\to{\mathbb{C}}$ be a Hermitian perfect pairing. Let $\mathcal{E}$ be the holomorphic bundle on $\Delta$ which is the Deligne extension of the flat bundle $\mathbb{L}\otimes{\mathcal{O}}_{\Delta^{\times}}$ [Del70, Prop. II.5.2]. Let $N\colon\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the residue of the flat logarithmic connection $\nabla\colon\mathcal{E}\to\frac{1}{z}\Omega^{1}_{X}(\mathcal{E})$ on $\Delta$. By continuous extension, $Q$ induces a pairing ׯ$\mathcal{E}\times\overline{\mathcal{E}}\to\mathcal{C}^{\infty}$ which restricts over the point $0\in\Delta$ to the canonical Hermitian pairing $Q_{0}\colon\mathcal{E}_{0}\times\overline{\mathcal{E}}_{0}\to{\mathbb{C}}$. Note that our $N$ differs from the one in [SS22] by a sign.
@@
    $\mathsf{F}$ continuously extends to $\mathcal{E}$ and its fibre over 00 is a mixed Hodge structure polarized by $N$ and $Q_{0}$.
@@
1) $\Rightarrow$ 2) is due to Schmid, see [CK89, (2.1)].
@@ 5.4. Reminder on Simpson’s non-abelian Hodge locus
Let $f\colon X\to S$ be a good holomorphic map. As recalled in Subsection 2.2, $f$ is a topological fibration, so the sheaf of pointed sets $R^{1}f_{*}{\mathrm{GL}}_{r}({\mathbb{C}})$ is a local system on $S$, with fibre $H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$, the set of isomorphism classes of rank $r$ complex local systems on $X_{s}$. Its associated étale space is a covering $\epsilon:T\to S$ ([Bou16, I.86 Prop. 9]). We endow $T$ with the usual complex manifold structure. As a topolocial space $T$ can be constructed as
@@ 6.1. Setting
over $X_{\Delta}$, where $\Delta\subset S$ is a suitable contractible open neighborhood of $s$. Let $\mathbb{L}_{\Delta}$ be the extension of the local system $\mathbb{L}$ to $X_{\Delta}$, let similarly $Q_{\Delta}$ denote the extension of the polarization. As in Appendix B the set of $t\in\Delta$ such that $(\mathbb{L}_{\Delta}|_{X_{t}},\mathsf{F}|_{X_{t}},Q_{\Delta}|_{X_{t}})$ is a -${\mathbb{C}}\text{-}{\sf PVHS}$ is open in $\Delta$, here we need the properness of $f$. By the same reasoning as in the proof of Theorem 3.1 the connected component of the non-abelian Hodge locus $\mathrm{NL}(f,r)$ containing $[\mathbb{L}_{s}]$, see Subsection 5.4, is a finite étale covering of $S$. So the monodromy orbit of $[\mathbb{L}_{s}]$ is finite.
@@ Appendix A Non-abelian Hodge loci (after Simpson)
By Theorem 5.1(iv) we have to show that for each simple constituent $\mathbb{M}$ of $\mathbb{L}$ the set of $s\in S$ such that $\mathbb{M}_{s}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ is closed analytic in $S$.
We fix such a simple constituent $\mathbb{M}$ and a Hermitian perfect pairing $Q\colon\mathbb{M}\times\overline{\mathbb{M}}\to{\mathbb{C}}$, which is unique up to a factor in $\mathbb{R}^{\times}$. In view of Theorem 5.1(ii) there are only finitely many numbers $h^{i}\geq 0$ ($i\in{\mathbb{Z}}$) which can occur up to shift as Hodge numbers of a -${\mathbb{C}}\text{-}{\sf PVHS}$ on $\mathbb{M}_{s}$ for some $s\in S$. In the following we fix such numbers $h^{i}$. For each component $E$ of $\overline{X}\setminus X$ we also fix numbers $h^{i}_{E,j}\geq 0$ for $|j|\leq r$ such that $h^{i}=\sum_{j}h^{i}_{E,j}$ for all $i\in{\mathbb{Z}}$. Clearly, also here there are only finitely many choices. Let $\mathrm{NL}_{\mathbb{M}}$ be the set of $s\in S$ such that $\mathbb{M}_{s}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ with Hodge numbers $h^{i}$ and polarization $Q|_{X_{s}}$ and such that the $j$-th monodromy graded piece $\mathrm{gr}^{\mathsf{M}}_{j}$ of $\mathbb{M}_{s}[E]$ has Hodge numbers $h^{i}_{E,j}$ for all components $E$ of $\overline{X}\setminus X$ and $j\in\mathbb{Z}$. Here we use the notation for the monodromy filtration as in [Del80, 1.7.8].
We have now reduced the proof of Theorem 5.10 to the construction of a proper holomorphic map of analytic spaces $\psi^{\circ}\colon G^{\circ}\to S$ such that $\mathrm{NL}_{\mathbb{M}}=\operatorname{{\rm im}}(\psi^{\circ})$, since then we conclude by Remmert’s proper mapping theorem. Let $\mathcal{E}$ be the Deligne extension of $\mathbb{M}\otimes_{{\mathbb{C}}}{\mathcal{O}}_{X}$ to a holomorphic bundle on $\overline{X}$.
Let $\mathrm{Gr}\to S$ be the projective map of analytic spaces, where $\mathrm{Gr}$ is the Grassmannian a point of which consists of $s\in S$ together with a filtration by holomorphic subbundles $\mathsf{F}^{i}\subset\mathcal{E}_{s}$ with $\mathrm{rank}(\mathsf{F}^{i}/\mathsf{F}^{i+1})=h^{i}$ for all $i\in{\mathbb{Z}}$. Let $G\hookrightarrow\mathrm{Gr}$ be the closed analytic subset consisting of those filtrations on fibres of $f$ which satisfy Griffiths transversality $\nabla\mathsf{F}^{i}\subset\Omega_{X_{s}}^{1}(\mathsf{F}^{i-1})$ and such that for each point $\{x\}=E_{s}$ at infinity and for each $j\in{\mathbb{Z}}$ the Hodge numbers of the filtrations induced on the monodromy graded pieces satisfy $\mathrm{rank}(\mathrm{gr}^{i}_{\mathsf{F}}{\rm gr}^{\mathsf{M}}_{j}\mathcal{E}_{x})=h^{i}_{E,j}$ for all $i$ and $j$.
@@
The set $G^{\circ}$ consisting of pairs $(s,\mathsf{F})\in G$ which define a polarizable variation of $\mathbb{C}$-Hodge structure on $X_{s}$ is open in $G$.
@@
Note that the points $(s,\mathsf{F})$ of $G$ in the claim correspond to those filtrations $\mathsf{F}$ such that there exists a perfect Hermitian pairing $Q_{s}$ for which the Hodge-Riemann relations are satisfied at every point of $X_{s}$, see Subsection 5.1. For a variation of complex Hodge structure the Hodge-Riemann relations are an open condition at every point of $X$, so the claim is easy in case $f$ is proper as then an open neighborhood of a fibre $X_{s}$ of $f$ contains the preimage of an open neighborhood of $s$ in $S$.
When $f$ is not proper the argument is more complicated. Consider a sequence $(s{(n)},\mathsf{F}{(n)})\in G$ converging to $(s,\mathsf{F})\in G^{\circ}$. In the sequel we fix local coordinates for the divisors at infinity on $\overline{X}$ in order to be able to speak about the limiting Hodge structure. Then by Proposition 5.3 we see that $\mathsf{F}_{x}$ for $\{x\}=E_{s}$ is the limiting mixed Hodge structure polarized by the induced pairing $Q_{x}\colon\mathcal{E}_{x}\times\overline{\mathcal{E}}_{x}\to{\mathbb{C}}$ and the nilpotent endomorphism $N_{x}$ on $\mathcal{E}_{x}$.
So by the openness of the mixed Hodge locus in the Grassmannian for $n\gg 0$ also $\mathsf{F}{(n)}_{E_{s{(n)}}}$ is a polarized mixed Hodge structure with respect to the polarization induced by $N$ and $Q$. Then by Proposition 5.3 and Remark 5.4 there is an open neighborhood $U_{x}$ of $x\in\overline{X}$ such that $\mathsf{F}{(n)}$ is a polarizable pure variation of Hodge structure on $U_{x}\cap X_{s{(n)}}$ for $n\gg 0$.
A similar open neighborhood $U_{x}$ of $x\in X$ exists for each $x\in X_{s}$ as explained at the beginning of the proof. By compactness of $\overline{X}_{s}$ there is finite union $U$ of such $U_{x}$ with $\overline{X}_{s}\subset U$. Then for $n\gg 0$ we obtain that $\overline{X}_{s{(n)}}\subset U$ and that $\mathsf{F}{(n)}$ is a polarizable pure Hodge structure at each point of $X_{s{(n)}}$. ∎
@@
The map $G^{\circ}\to S$ is proper.
In the proof of Claim A.3 we need the following continuity result for a fibrewise pluri-harmonic metric due to T. Mochizuki [Moc09, Prop. 4.2]. We note that $\det\mathbb{M}$ is finite by Theorem 5.1(v), so we can fix a positive flat Hermitian metric $h^{\det}$ on $\det\mathbb{M}$, which is automatically pluri-harmonic.
@@
Let $h\colon(\mathbb{M}\otimes\mathit{C}^{\infty}_{X})\times(\overline{{\mathbb{M}}\otimes\mathit{C}^{\infty}_{X}})\to\mathit{C}^{\infty}_{X}$ be the unique fibrewise pluri-harmonic, tame and purely imaginary metric with determinant $h^{\det}$. Then $h$ is continuous.
@@
Consider a sequence $(s{(n)},\mathsf{F}{(n)})\in G^{\circ}$ such that $s{(n)}$ converges to $s\in S$. We have to show that a subsequence converges in $G^{\circ}$. By the properness of the Grassmannian we can assume that $(s{(n)},\mathsf{F}{(n)})$ converges to $(s,\mathsf{F})\in G$. We have to prove that $(s,\mathsf{F})\in G^{\circ}$, i.e. we have to check the Hodge-Riemann relations at all points $x\in X_{s}$. Note that the orthogonal complement $(\mathsf{F}^{i}_{x})^{\perp}$ of $\mathsf{F}^{i}_{x}$ in $\mathbb{L}_{x}$ with respect to $Q_{x}$ and with respect to the fibrewise pluri-harmonic metric $h$ from Proposition A.4 coincide by continuity since they coincide for $\mathsf{F}(n)^{i}$ by the description of the metric $h$ as a Hodge metric in terms of $Q$ on the fibre over $s{(n)}$. As $h$ is positive definite we deduce that $\mathbb{L}_{x}=\mathsf{F}^{i}_{x}\oplus(\mathsf{F}^{i}_{x})^{\perp}$. In other words $Q_{x}$ restricted to $\mathsf{F}^{i}$ is a perfect pairing. For the same reason $(-1)^{i}Q_{x}$ restricted to $\mathsf{F}^{i}_{x}\cap(\mathsf{F}^{i+1}_{x})^{\perp}$ is positive definite as by continuity it agrees with $h_{x}$ there up to a positive factor. ∎

### v4

@@ 1. Introduction
    the filtration ${\mathsf{F}}_{s}$ extends to a (relative) Griffiths transversal filtration by subbundles $\hat{\mathsf{F}}$ of the formal isomonodromic Deligne extension (^,^:^$(\hat{\mathcal{E}},\hat{\nabla}\colon\hat{\mathcal{E}}\to\hat{\Omega}^{1}_{\hat{\overline{X}}/\hat{S}}(\log D)(\hat{\mathcal{E}}))$.
Here “essentially extends” means that it extends after base change with respect to a finite, étale, surjective covering of $S$, see Subsection 3.1 for a precise formulation. In 5) $\hat{S}$ is the formal scheme of $S$ along $s$ and $\hat{\overline{X}}$ the formal scheme of $\overline{X}$ along $\overline{X}_{s}$. The formal isomonodromic Deligne extension $(\hat{\mathcal{E}},\hat{\nabla})$ is the formal completion of the log flat bundle on $\overline{X}_{\Delta}$ associated by Deligne [Del70, Prop. 5.4, Rmk. 5.5] to the local system on $X_{\Delta}$ which canonically extends $\mathbb{L}_{s}$. Recall that the filtration ${\mathsf{F}}_{s}$ canonically extends to a filtration by subbundles of the Deligne extension over $\overline{X}_{s}$ by Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], so part 5) can be interpreted as a formal deformation problem for the latter filtration of bundles from $\overline{X}_{s}$ to $\hat{\overline{X}}$.
@@
Acknowledgment. It is a pleasure to thank Alexander Petrov for numerous enlightening discussions and for sharing his insights into isomonodromic deformations of Hodge structures. In particular, he alerted us on a gap in a previous version of Appendix A. We thank Claude Sabbah for kindly helping us with references on variations of Hodge structure. We thank Daniel Litt for various helpful discussions on local systems in particular on the condition 5) in Theorem 1.1. We are grateful to Takuro Mochizuki for explaining to us what is known and what can be expected about families of pluri-harmonic metrics. We thank Xinwen Zhu for an enlightening discussion on the analogy between Theorem 1.1 and his own theorem on the preservation of the de Rham property of a $p$-adic local system, see [LZ17].
@@ 2.2. Geometric examples
Let $X$ and $S$ be connected complex manifolds and let $f\colon X\to S$ be a good holomorphic map as defined in Section 1. Fix a point $x\in X$ and set $s=f(x)$. By Thom’s first isotopy lemma [GM88, Thm. I.1.5] or Ehresmann’s theorem in the proper case, $f$ is a topological fibre bundle. We obtain an exact sequence of homotopy groups
@@ 3.1. Integral Fixed Part Theorem
As $f$ is a topological fibre bundle, $\pi_{1}(S,s)$ acts on the cohomology $H^{i}(X_{s},{\mathbb{Q}})$. We denote by $\widehat{S}$ the formal completion of $S$ along $s$ and by $\widehat{X}$ the formal completion of $X$ along $X_{s}$.
@@ 5.1. Polarized variations of ${\mathbb{C}}$-Hodge structure
Let us recall for later reference one of the main results about degeneration of pure ${\mathbb{C}}$-Hodge structure. Let $\mathbb{L}$ be a unipotent complex local system on $\Delta^{\times}=\Delta\setminus\{0\}$ with $\Delta=\{z\in{\mathbb{C}}\,|\,|z|<1\}$. Let $Q\colon\mathbb{L}\times\overline{\mathbb{L}}\to{\mathbb{C}}$ be a Hermitian perfect pairing. Let $\mathcal{E}$ be the holomorphic bundle on $\Delta$ which is the Deligne extension of the flat bundle $\mathbb{L}\otimes{\mathcal{O}}_{\Delta^{\times}}$ [Del70, Prop. II.5.2]. Let $N\colon\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the residue of the flat logarithmic connection $\nabla\colon\mathcal{E}\to\frac{1}{z}\Omega^{1}_{X}(\mathcal{E})$ on $\Delta$. One can construct an induced Hermitian pairing $Q_{0}\colon\mathcal{E}_{0}\times\overline{\mathcal{E}}_{0}\to{\mathbb{C}}$ using a (tangential) base point. Note that our $N$ differs from the one in [SS22] by a sign.
@@
    $\mathsf{F}$ extends to a filtration by holomorphic subbundles of $\mathcal{E}$ and its fibre over 00 induces a mixed Hodge structure polarized by $N$ and $Q_{0}$, i.e. a nilpotent orbit by [CK89, (2.3)].
@@
1) $\Rightarrow$ 2) is Schmid’s nilpotent orbit theorem, see [CK89, (2.1)].
@@
Let now $X$ be a Riemann surface with compactification $\overline{X}$ and let $(\mathbb{L},\mathsf{F},Q)$ be a -${\mathbb{C}}\text{-}{\sf PVHS}$ on $X$. Assume that the local monodromies at infinity of $\mathbb{L}$ are unipotent. Then the anti-holomorphic Deligne extension $\overline{\mathcal{E}}$ of the complex conjugate local system ${\overline{\mathbb{L}}}$ is the complex conjugate bundle of $\mathcal{E}$. By [EV86, App.B] ${\rm deg}(\overline{\mathcal{E})}={\rm deg}(\mathcal{E})=0$. By the unipotence of local monodromy the formation of the Deligne extension is compatible with the dual bundle.
By Schmid’s nilpotent orbit theorem $\mathsf{F}^{a}$ extends to a holomorphic subbundles of $\mathcal{E}$ and its orthogonal complement $\overline{\mathsf{F}}^{a}=(\mathsf{F}^{-a})^{\perp}$ extends to an anti-holomorphic subbundle of $\overline{E}$, which we denote by the same symbols. So $\overline{\mathsf{F}}^{a}$ is the complex conjugate dual of $\mathcal{E}/{\mathsf{F}}^{-a}$. We conclude
Lemma 5.5.
For all integers $a$ it holds
$\deg({\mathsf{F}}^{a})+\deg(\overline{\mathsf{F}}^{-a})=0.$
@@ 5.4. Reminder on Simpson’s non-abelian Hodge locus
Let $f\colon X\to S$ be a good holomorphic map. As recalled in Subsection 2.2, $f$ is a topological fibre bundle, so the sheaf of pointed sets $R^{1}f_{*}{\mathrm{GL}}_{r}({\mathbb{C}})$ is a local system on $S$, with fibre $H^{1}(X_{s},{\mathrm{GL}}_{r}({\mathbb{C}}))$, the set of isomorphism classes of rank $r$ complex local systems on $X_{s}$. Its associated étale space is a covering $\epsilon:T\to S$ ([Bou16, I.86 Prop. 9]). We endow $T$ with the usual complex manifold structure. As a topolocial space $T$ can be constructed as
@@ 6.1. Setting
over $X_{\Delta}$, where $\Delta\subset S$ is a suitable contractible open neighborhood of $s$. Let $\mathbb{L}_{\Delta}$ be the extension of the local system $\mathbb{L}$ to $X_{\Delta}$, let similarly $Q_{\Delta}$ denote the extension of the polarization. As in Appendix B, see Claim A.3, the set of $t\in\Delta$ such that $(\mathbb{L}_{\Delta}|_{X_{t}},\mathsf{F}|_{X_{t}},Q_{\Delta}|_{X_{t}})$ is a -${\mathbb{C}}\text{-}{\sf PVHS}$ is open in $\Delta$, note that in the non-proper case this observation relies on [KU09, Prop. 7.1.1]. By the same reasoning as in the proof of Theorem 3.1 the connected component of the non-abelian Hodge locus $\mathrm{NL}(f,r)$ containing $[\mathbb{L}_{s}]$, see Subsection 5.4, is a finite étale covering of $S$. So the monodromy orbit of $[\mathbb{L}_{s}]$ is finite.
@@ Appendix A Non-abelian Hodge loci (after Simpson)
By Theorem 5.1(iv) we have to show that for each simple constituent $\mathbb{M}$ of $\mathbb{L}$ the set of $s\in S$ such that $\mathbb{M}_{s}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ is closed analytic in $S$. We will see that this holds as a consequence of the key Proposition A.2 below.
We fix such a simple constituent $\mathbb{M}$ and a Hermitian perfect pairing $Q\colon\mathbb{M}\times\overline{\mathbb{M}}\to{\mathbb{C}}$, which is unique up to a factor in $\mathbb{R}^{\times}$. Let $\mathcal{E}$ be the Deligne extension of $\mathbb{M}\otimes_{{\mathbb{C}}}{\mathcal{O}}_{X}$ to a holomorphic bundle on $\overline{X}$.
Grothendieck constructed a certain flag scheme, see [HL10, 2.A.1], which for us is a morphism of analytic spaces $\psi\colon G\to S$, where a point $\xi\in G$ corresponds to a descending finite filtration by coherent subsheaves $\mathsf{F}^{i}_{\xi}$ ($i\geq 0$) of $\mathcal{E}_{\psi(\xi)}$ such that $\mathcal{E}_{\psi(\xi)}/\mathsf{F}^{i}_{\xi}$ is torsion free for all $i$ and such that Griffiths transversality $\nabla\mathsf{F}^{i}\subset\Omega_{X_{\psi(\xi)}}^{1}(\mathsf{F}^{i-1})$ holds for all $i\in\mathbb{Z}$. In order to get rid of the ambiguity of shifting the filtration, we also assume that we consider only such filtrations with $\mathsf{F}^{0}_{\xi}=\mathcal{E}_{\psi(\xi)}$ and $\mathsf{F}^{1}_{\xi}\neq\mathcal{E}_{\psi(\xi)}$. Let $G^{0}$ be the subset of $\xi\in G$ corresponding to the filtrations by subbundles such that $\mathsf{F}_{\xi}|_{X_{{\psi(\xi)}}}$ is a variation of Hodge structure polarized by $Q$.
Proposition A.2.
The following holds:
  • 1)
    $G$ is the countable disjoint union of analytic spaces which are projective over $S$, more precisely the points $\xi\in G$ such that the ranks and degrees of all coherent sheaves ${\rm gr}_{\mathsf{F}}^{i}\mathcal{E}_{\psi(\xi)}$ are fixed form a clopen subset of $G$ which is projective over $S$;
  • 2)
    $G^{0}$ consists of finitely many connected components of $G$.
Proposition A.2 and Remmert’s proper mapping theorem imply that $\psi(G^{0})$ is a closed analytic subset of $S$. On the other hand by analytic continuation of the Hodge filtration along the Deligne extension, i.e. the nilpotent orbit theorem [CK89, (2.1)], this is precisely the above set of $s\in S$ such that $\mathbb{M}_{s}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ polarized by $Q$.
Proof of Proposition A.2.
Part 1) is immediate from Grothendieck’s results, see [HL10, 2.A.1].
For part 2) we have to show that there are only finitely many collections of numbers which can occur as the ranks and degrees as in 1) for elements of $G^{0}$. The finiteness of the number of ranks follows from the fact that we assumed $\mathsf{F}_{\xi}^{0}\not\subset\mathsf{F}^{1}_{\xi}\neq\mathcal{E}_{\psi(\xi)}$ and that there can be no gap in the Hodge filtration of a simple local system by Theorem 5.1(ii), i.e. if the rank of ${\rm gr}_{\mathsf{F}}^{i}\mathcal{E}_{\psi(\xi)}$ vanishes for some $i>0$ then it also vanishes for all larger indices $i$. Similarly, the degrees are bounded as the associated log Higgs bundle is stable [Sim90, Thm.5] and we can therefore apply [Nit91, Prop. 3.2] to $L_{s}=\omega_{\bar{X}_{s}}(\bar{X}_{s}\setminus X_{s}),\ s\in S$ which has a constant degree for all fibers $X_{s}$ of $X\to S$.
To conclude the proof of part 2) we have to show that $G^{0}$ is both open and closed in $G$, which is a consequence of Claim A.3 and Claim A.4. ∎
@@
The subset $G^{0}$ of $G$ is open.
@@
Let us consider the subset $Z\subset G\times_{S}\overline{X}$ consisting of those points $(\xi,x)$ such that either $x\in X$ and $\mathsf{F}_{\xi,x}$ is a Hodge filtration polarized by $Q$ or $x\in\overline{X}\setminus X$ and $\mathsf{F}_{\xi,x}$ together with the nilpotent monodromy operator and $Q$ is a nilpotent orbit. We have to show that $Z$ is an open subset. Indeed this implies that $G^{0}$ is open as it is then equal to the complement of the image of the closed subset $(G\times_{S}\overline{X})\setminus Z$ along the proper projection $G\times_{S}\overline{X}\to G$.
The subset $Z$ is open around $G\times_{S}X$, because the Hodge-Riemann inequality is an open condition. Around $G\times_{S}(\overline{X}\setminus X)$ the subset $Z$ is open as the property of being a nilpotent orbit, granted Griffiths transversality holds, is an open condition by [KU09, Prop. 7.1.1] and as a Griffiths transversal filtration by subbundles, which can be approximated by a nilpotent orbit, is a variation of Hodge structure locally by Remark 5.4. ∎
@@
The subset $G^{\circ}$ of $G$ is closed.
In the proof of Claim A.4 we need the following continuity result, Proposition A.5, for a fibrewise harmonic metric due to T. Mochizuki [Moc09, Prop. 4.2]. We note that $\det\mathbb{M}$ is finite by Theorem 5.1(v), so we can fix a positive flat Hermitian metric $h^{\det}$ on $\det\mathbb{M}$, which is automatically pluri-harmonic, to be concrete take $h^{\det}=\pm Q^{\det}$.
@@
Let $h\colon\mathcal{E}|_{X}\times\overline{\mathcal{E}}|_{X}\to\mathbb{C}$ be the unique fibrewise harmonic, tame and purely imaginary metric with determinant $h^{\det}$. Then $h$ is continuous.
@@
Consider a sequence $\xi(n)\in G^{\circ}$ converging to $\xi\in G$. In order to show that $\xi\in G^{0}$ we have to verify two properties: 1) $\mathsf{F}_{\xi}^{i}$ is a subbundle of $\mathcal{E}_{\psi(\xi)}$ for all $i$, 2) the $\mathsf{F}_{\xi}|_{X_{\psi(\xi)}}$ form a variation of Hodge structure polarized by $Q$.
Consider a point $x\in X_{\psi(\xi)}$ around which 1) holds; note that a priori there might be finitely many points for which it fails. The orthogonal complement $(\mathsf{F}^{i}_{\xi,x})^{\perp}$ of $\mathsf{F}^{i}_{\xi,x}$ in $\mathbb{M}_{x}$ with respect to $Q_{x}$ and with respect to the fibrewise harmonic metric $h_{x}$ from Proposition A.5 coincide by continuity since they coincide on $\mathcal{E}_{\psi(\xi(n))}^{i}$ for all $n$ by the description of the metric $h$ as a Hodge metric in terms of $Q$ on the fibre of $X$ over $\psi(\xi(n))$. As $h$ is positive definite we deduce that $\mathbb{M}_{x}=\mathsf{F}^{i}_{\xi,x}\oplus(\mathsf{F}^{i}_{\xi,x})^{\perp}$. In other words $Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}$ is a perfect pairing. For the same reason $(-1)^{i}Q_{x}$ restricted to $\mathsf{F}^{i}_{\xi,x}\cap(\mathsf{F}^{i+1}_{\xi,x})^{\perp}$ is positive definite as by continuity it agrees with $h_{x}$ there. So around $x$ we get a variation of Hodge structure $\mathsf{F}_{\xi}$.
In order to conclude we have to show that the torsion vanishes in the coherent sheaf $\mathcal{E}_{\psi(\xi)}/\mathsf{F}^{i}_{\xi,\psi(\xi)}$ for any $i$. Let $\overline{\xi}(n)$ correspond to the orthogonal Hodge filtration $\overline{\mathsf{F}}^{a}=({\mathsf{F}}^{-a})^{\perp}$ on $X_{\psi(\xi(n))}$ continuously extended to the anti-holomorphic Deligne extension ¯$\overline{\mathcal{E}}_{\psi(\xi(n))}$ on $\overline{X}_{\psi(\xi(n))}$. By the properness of $G$ over $S$, Proposition A.2 1), we can replace $(\xi(n))_{n}$ by a subsequence such that $\overline{\xi}(n)$ converges to some $\overline{\xi}$ in the anti-holomorphic version of $G$. By the continuity of the degree on $G$ and by Lemma 5.5 we obtain
$\deg(\mathsf{F}^{i}_{\xi,\psi(\xi)})+\deg(\mathsf{F}^{-i}_{\overline{\xi},\psi(\xi)})=0$
for all $i$ as this holds for the Hodge filtrations corresponding to $\xi(n)$ for all $n$. By Schmid’s nilpotent orbit theorem, see [CK89, (2.1)], the saturations $\tilde{\mathsf{F}}_{\xi,\psi(\xi)}$ of $\mathsf{F}_{\xi,\psi(\xi)}$ and $\tilde{{\mathsf{F}}}_{\overline{\xi},\psi(\xi)}$ of ${\mathsf{F}}_{\overline{\xi},\psi(\xi)}$ form a variation of Hodge structure on $X_{\psi(\xi)}$ of weight $-1$, so
$\deg(\tilde{\mathsf{F}}^{i}_{\xi,\psi(\xi)})+\deg(\tilde{{\mathsf{F}}}^{-i}_{\overline{\xi},\psi(\xi)})=0$
by Lemma 5.5. Therefore, the inequalities
$\deg(\mathsf{F}^{i}_{\xi,\psi(\xi)})\leq\deg(\tilde{\mathsf{F}}^{i}_{\xi,\psi(\xi)}),\quad\deg({\mathsf{F}}^{-i}_{\overline{\xi},\psi(\xi)})\leq\deg(\tilde{{\mathsf{F}}}^{-i}_{\overline{\xi},\psi(\xi)})$
have to be equalities and consequently $\mathsf{F}^{i}_{\xi,\psi(\xi)}=\tilde{\mathsf{F}}^{i}_{\xi,\psi(\xi)}$. ∎